
None.
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2014-04-14 | 9.0 | 1.5 | 0.0 | 0.0 | 0.0 | 10.5 |
Links
[1] http://www.ctio.noao.edu/noao/field-collection/field-nr-conditions/1488/delete?destination=print/node/5417
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-conditions/1488/edit?destination=print/node/5417
[3] mailto:jelias@noao.edu
[4] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/1489/delete?destination=print/node/5417
[5] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/1489/edit?destination=print/node/5417
[6] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/5417?destination=print/node/5417
Another successful night for COSMOS comissioning. We observerd Hamuy spectrophotometric standards, some PNe for PR, and and astrometric field.
We also tested the functionality of the Cass ADC. We will compare these rusults to ights in which we did not use the ADC.
At the end of the night, we have closed the COSMOS observing interface, powered doen the IEB, and turned off the Torrent controller. We have also removed
the Sloan 4"x4" filter set from COSMOS and returned the filters to the proper location on floor 2. Filters on loan from SOAR will be sent back to Pachon tomorrow.
COSMOS equiment and tools left in their room on the P floor.